$ 4.800 5,250 5,300 3,000 30,300 Accounts eceivable Prepaid advertising. .. . Trucks._.. . $10,000 Equipment . . . . Accumulated depreciation-Equipment... Accounts payable... Uneaned service fees . . 7,600 2,100 6.700 10,000 16,050 Retained eamings Service fees eaned... Rent expense.. 7.500 80,500 29,800 10,200 4,900 8,000

Prepare a detailed Balance Sheet and also calculate the following ratios:

A) Current Ratio

B) Working Capital

C) Assets Turnover

D) Fixed Assets Turnover

E) A/R Turnover

F) Debt Ratio
Answer #1

Solution:

A. Current Ratio = Current assets / Current liabilities

   = (4,800 + 5,250 +5,300 + 3,000) / (1,200 +6,700) = 2.32 (Rounded off )

B.Working capital = Current assets - Current liabilities

   = (4,800 + 5,250 +5,300 + 3,000) - (1,200 +6,700) = 10,450

C. Assets Turnover = Net sales / Total assets

= 80,500 / 44,150 = 1.82 (Rounded off)

D. Fixed assets turnover = Net sales / Total Fixed assets

= 80,500 / 25,800 = 3.12 (Rounded off)

E. A/R turnover = Net credit sales / Average Accounts Receivable

   = 80,500 / 5,250    (Here Service is assumed to be credit)

   = 15.33 (Rounded off)

F. Debt ratio = Total debt / Total assets

= 1,200 / 44,150 = 0.027